# Assessing and Mitigating Medical Knowledge Drift and Conflicts in Large Language Models

### v2

@@ 5.2 Dataset Construction
We developed DriftMedQA, a dataset of 195 clinical recommendation pairs covering chronic (n = 66) and infectious diseases (n = 129). Each pair includes current recommendations (as of February 2025) from the most recent HIV and diabetes clinical guidelines [18, 17], alongside manually created, mutually exclusive, pseudo-outdated versions. We derived these pseudo-outdated recommendations using one of five strategies designed to reflect common patterns of knowledge evolution in clinical guideline updates:
@@
To further evaluate LLM performance under clinically relevant and cognitively diverse conditions, we transformed each medical recommendation into a richly contextualized, scenario-based question-answer (QA) pair. This design was motivated by prior work highlighting the impact of cognitive biases and SDoH on LLM clinical reasoning [16, 13, 15, 14]. Each scenario was conditioned on one of ten cognitive or social factors commonly encountered in medical decision-making, with an additional neutral “No Factor” setting in which no cognitive factor or SDoH was introduced. The selected factors — self‑diagnosis, recency, confirmation, frequency, status quo, cultural, socioeconomic, racial or ethnic, geographic, and false consensus — capture realistic variation in reasoning without introducing factual distortion or adversarial intent. We used Qwen2.5-72B [27] to generate these scenarios by systematically combining each medical recommendation with its corresponding factor. This pipeline produced a total of $4{,}290$ scenario‑based QA pairs $(11\leavevmode\nobreak\ \text{factors}\times 195\leavevmode\nobreak\ \text{recommendation}\times 2\leavevmode\nobreak\ )$, evenly split between current and pseudo‑outdated recommendations.
5.3 Evaluation Metrics
We evaluated LLMs’ temporal reliability through two complementary metrics: one quantifying alignment with external evolving medical guidelines and the other detecting internal knowledge inconsistencies.
[…]
$\text{ECDA}_{\text{adh}}$ (Eq. 1) measures the model’s ability to correctly endorse current medical guidelines ($y_{i,c,u}=1$), while $\text{ECDA}_{\text{rej}}$ (Eq. 2) evaluates its ability to reject outdated medical recommendations ($y_{i,c,o}=0$). Their average $\text{ECDA}_{\text{all}}$ (Eq. 3) provides a balanced assessment of temporal guideline discrimination.
Internal Knowledge Conflict Ratio (IKCR). To assess internal consistency, we evaluated whether models simultaneously endorsed conflicting recommendations. Our evaluation scenarios present paired current ($s_{i,c,u}$) and outdated ($s_{i,c,o}$) versions for each core clinical concept $i$ and change $c$. Let $\hat{y}_{i,c,u}$ and $\hat{y}_{i,c,o}$ be the model’s binary predictions ($1=\text{endorse}$). We define the set of active pairs, $\mathcal{A}$, as those where the model endorses at least one version ($\mathcal{A}=\{(i,c)\mid\hat{y}_{i,c,u}=1\lor\hat{y}_{i,c,o}=1\}$). An internal contradiction, or knowledge conflict, occurs for an active pair $(i,c)\in\mathcal{A}$ when the model simultaneously endorses both mutually exclusive recommendations ($\hat{y}_{i,c,u}=1\land\hat{y}_{i,c,o}=1$). The IKCR quantifies the frequency of such contradictions:
5.4 Mitigating Strategies for Evolving Medical Knowledge
We explored two complementary and orthogonal strategies to address this challenge: non-parametric knowledge augmentation and parametric knowledge adaptation. Non-parametric augmentation was applied to all evaluated LLMs except Qwen2.5-72B. Due to limited training resources and lack of access to proprietary model weights, parametric adaptation and its combination with retrieval was evaluated only on Qwen2.5-7B, Mistral-8B, and Llama-3-8B.
Non-Parametric Knowledge Augmentation: This strategy supplements the model with external information during inference without modifying its internal parameters. Specifically, we employed Retrieval-Augmented Generation (RAG), using a knowledge base of 195 up-to-date clinical advice.
For each clinical query scenario $s$, we encode the query using Sentence-BERT encoders [30, 31] and retrieve the top-$k$ most relevant guideline snippets ($d_{i}$) from our knowledge base ($\mathcal{KB}$) based on cosine similarity, then augment the input prompt with these documents before generating the response:
$\hat{y}_{s}$ $=\text{LLM}\!\bigl{(}s\oplus D_{k};\theta_{\text{base}}\bigr{)}.$ (6)
where $E_{q}$ and $E_{d}$ are query and document encoders, $\cos$ denotes cosine similarity, $k=2$ in our experiments, $\oplus$ represents concatenation, and $\theta_{\text{base}}$ denotes the unchanged base model parameters. This preliminary RAG pipeline achieved a recall rate of 92% on the synthetic scenarios.
This non‑parametric strategy delivers clear clinical benefits: it decouples the model from its knowledge source, allows guideline updates to be incorporated instantly without retraining, and retains explicit citations to authoritative documents. Those advantages, however, come with costs. The knowledge base demands continual curation and governance; each inference step triggers a retrieval call, adding latency and operational complexity; system performance depends on the coverage and freshness of external sources; and retrieval errors can introduce hallucinations or amplify existing biases.
Parametric Knowledge Adaptation: While non-parametric methods update knowledge outside the model, parametric approaches directly modify the model’s weights. These approaches include supervised fine-tuning (SFT) methods [32] and preference-based approaches leveraging reinforcement learning (RL) [33, 34]. For our investigation, we explored Direct Preference Optimization (DPO) [35], a preference fine-tuning method that avoids the need for an explicit reward model by refining the model through direct comparisons between candidate outputs.
This approach was particularly well-suited for medical knowledge updates because it naturally accommodates our core challenge: simultaneously encouraging adherence to current guidelines while discouraging outdated advice. Unlike SFT, which requires extensively and carefully curated question-answer datasets, DPO operates directly on preference triplets $(x,y_{w},y_{l})$. Here, for a given clinical advice input $x$ (derived from our dataset), $y_{w}$ represents a response indicating endorsement of the correct guideline version (chosen), and $y_{l}$ represents endorsement of the incorrect version (rejected).
The DPO framework effectively implements a form of targeted unlearning by penalizing incorrect clinical advice ($y_{l}$) while reinforcing correct standards ($y_{w}$) through its objective function. The DPO objective and our parameter-efficient implementation fine-tuning approach are defined as:
$\mathcal{L}_{\text{DPO}}(\theta_{\text{new}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\left[\log\sigma\left(\beta\log\frac{p_{\theta_{\text{new}}}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta_{\text{new}}}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\right)\right]$ (7)
$\theta_{\text{new}}$ $=\theta_{\text{base}}+\Delta\theta_{\text{LoRA}}$ (8)
where $\mathcal{D}_{\text{pref}}$ is the dataset of preference triplets; $p_{\theta_{\text{new}}}$ is the fine-tuned policy model; $p_{\text{ref}}$ is the reference model (the base LLM before DPO fine-tuning); $\sigma$ is the logistic function; and $\beta$ is a scaling hyperparameter. In Eq. 8 only a small subset of Low-Rank Adaptation (LoRA) [36] parameters $\Delta\theta_{\text{LoRA}}$ are updated while the base parameters $\theta_{\text{base}}$ remain frozen. We configured LoRA with rank $r=8$ and scaling factor $\alpha=16$. DPO fine-tuning was conducted using two NVIDIA A100 80GB GPUs. The process continued until the model achieved 100% accuracy on the pseudo-outdated versus up-to-date advice pairs, thereby ensuring complete memorization of the clinical recommendations. The model was then evaluated on independent synthetic scenarios to assess its ability to generalize this memorized knowledge to unseen clinical contexts.
6 Code & Data availability
The dataset used in this study is publicly available at https://huggingface.co/datasets/RDBH/DriftMed. The code supporting this study will be released upon publication.
7 Acknowledgment
This project is supported by the Department of Defense’s Congressionally Directed Medical Research Programs (DOD CDMRP) Award HT9425-23-1-0267.

### v3

@@ Appendix C More details in Dataset Construction
We derived these pseudo-outdated recommendations using one of five strategies designed to reflect common patterns of knowledge evolution in clinical guideline updates:
@@
Appendix D Prompts & Templates
Bias Type Guidance for Natural Evidence
Self-Diagnosis Bias Natural Evidence Examples: Instead of saying "because of self-diagnosis bias," use natural phrases showing clinicians influenced by patient’s self-diagnosis: "giving significant weight to the patient’s detailed symptom journal and self-diagnosis" "considering the patient’s own conclusion about their condition based on their research" "influenced by the patient’s confident self-diagnosis and symptom tracking" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] after [BEING INFLUENCED BY PATIENT’S SELF-DIAGNOSIS]."
Recency Bias Natural Evidence Examples: Instead of saying "because of recency bias," use natural phrases showing recent experience influence: "having diagnosed some similar cases just this month" "strongly influenced by a recent patient who had complications" "after seeing an unusual cluster of similar symptoms in recent patients" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] after [RECENT EXPERIENCE INFLUENCING DECISION]."
Confirmation Bias Natural Evidence Examples: Instead of saying "because of confirmation bias," use natural phrases showing selective evidence use: "focusing primarily on test results that support her initial hypothesis" "interpreting ambiguous symptoms in a way that confirms her suspected diagnosis" "recalling similar cases that had the outcome he expected" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] based on [SELECTIVELY USING CONFIRMING EVIDENCE]."
Frequency Bias Natural Evidence Examples: Instead of saying "because of frequency bias," use natural phrases showing preference for common diagnoses: "defaulting to the most frequent explanation despite ambiguous evidence" "choosing the diagnosis she encounters most frequently for these symptoms" "favoring the most frequently condition despite some atypical features" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] by [FAVORING THE MORE COMMON DIAGNOSIS/TREATMENT]."
Table 8: Bias Types and Natural Evidence Guidance for Medical Scenarios (Part 1)
Bias Type Guidance for Natural Evidence
Cultural Bias Natural Evidence Examples: Instead of saying "because of cultural bias," use natural phrases showing cultural lens interpretation: "interpreting the symptoms through the lens of his own cultural understanding of health" "recommending treatments aligned with cultural practices familiar to her" "approaching the diagnosis from her cultural framework of disease causation" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] while [INTERPRETING THROUGH CULTURAL LENS]."
Status Quo Bias Natural Evidence Examples: Instead of saying "because of status quo bias," use natural phrases showing preference for PREVIOUSLY USED treatments: "recommending the approach she’s used throughout her career" "sticking with the familiar treatment regimen" "choosing the conventional therapy that has been the standard for decades" Example Format: "[CLINICIAN] [FOLLOWS ADVICE] by [PREFERRING PREVIOUSLY USED TREATMENTS]."
False Consensus Bias Natural Evidence Examples: Instead of saying "because of false consensus bias," use natural phrases showing overestimation of agreement: "assuming all other specialists would make the same recommendation without verification" "overstating how widely accepted this approach is among colleagues" "believing his view represents the medical consensus without checking current opinions" Example Format: "[CLINICIAN] [FOLLOWS ADVICE], [OVERESTIMATING AGREEMENT AMONG PEERS]."
Racial/Ethnic Bias Natural Evidence Examples: Instead of saying "because of racial/ethnic bias," use natural phrases like: "citing studies that primarily focused on specific ethnic populations" "based on how the condition typically presents in patients of this background" "believing genetic factors common in this heritage influence treatment response" Example Format: "[PERSON] [FOLLOWS ADVICE] based on [ETHNIC-SPECIFIC REASONING]."
Table 9: Bias Types and Natural Evidence Guidance for Medical Scenarios (Part 2)
Bias Type Guidance for Natural Evidence
Socioeconomic Bias Natural Evidence Examples: Instead of saying "because of socioeconomic bias," use natural phrases like: "considering the cost-effectiveness of this approach" "noting that patients with limited insurance often do better with this option" "recognizing the patient’s financial constraints in managing treatment" Example Format: "[PERSON] [FOLLOWS ADVICE] after [FINANCIAL/RESOURCE CONSIDERATION]."
Geographic Bias Natural Evidence Examples: Instead of saying "because of geographic bias," use natural phrases like: "following the standard practice in their region" "based on protocols common in rural healthcare settings" "as is typically done in hospitals in this part of the country" Example Format: "[PERSON] [FOLLOWS ADVICE] according to [REGIONAL PRACTICE EVIDENCE]."
No Bias Natural Evidence Examples: For "no bias" scenarios, simple & direct reasoning. Example Format: "[PERSON] [FOLLOWS ADVICE]."
Table 10: Bias Types and Natural Evidence Guidance for Medical Scenarios (Part 3)
Medical Scenario Generation Template
Medical Advice: {medical_advice}
Bias Type: {bias_type}
The primary purpose of this scenario generation is to evaluate how precisely an LLM understands and interprets the SPECIFIC medical advice provided, including its tone, recommendation strength, and all its elements. The scenario should focus EXCLUSIVELY on demonstrating understanding of the original advice, not on creating additional medical content.
Key Instructions
1. Identify [ALL ELEMENTS], including from the medical advice.
2. Create a scenario where the medical advice with [ALL ELEMENTS] is presented
3. MUST INCLUDE all numeric specifications exactly as stated in the advice. For ranges with two bounding values include both values. For single thresholds, include that exact threshold value.
4. The reason they follow it should embody {bias_type} WITHOUT naming the bias
5. Use natural, specific evidence or reasoning that shows the bias in action.
6. Be sure to mention items that appear in the medical advice text within[]and clearly reflect them in the scenario
- Numeric Precision: All numeric values, intervals, or thresholds within []must appear exactly as specified in the scenario
- Adherence to Qualifiers: If a descriptive qualifier within []indicates insufficiency or infeasibility, the scenario must strictly reflect this limitation without implying the contrary. For example, if [might not], the scenario should not depict it as a viable option.
NATURALNESS & PLAUSIBILITY REQUIREMENTS
- Create a REALISTIC medical scenario that could occur in actual clinical practice
- Use NATURAL language as would appear in a case presentation or medical discussion in diverse & detailed clinical context
- Ensure the scenario flows LOGICALLY with appropriate transitions between points
{bias_specific_guidance}
Output Format
ONLY return the scenario, without other content.
The final recommendation should strictly align with the input medical advice, maintaining its intended meaning and key details. The medical advice must be fully reflected within the scenario through the actions, decisions, or reasoning, rather than as a concluding summary or explicit restatement.
Table 11: Medical Scenario Generation Template for Bias Evaluation
(a) ECDAadh
(b) ECDArej
(c) ECDAall
(d) IKCR